3. Garza Electronics expects to sell 500

units in January, 250 units in February,

and 1 .000 units in March. January's
beginning inventory is 700 units.
Expected sales for the whole year are
7,200 units. Garza has decided on a ]eve]
monthly production schedule of 600 units
(7,200 units/12 months : 600 units per
month). What is the expected end.of.
month inventory for January, February,
and March? Show the beginning
inventory, production, and sales for each
month to arrive at ending inventory.
Answer #1
Month

Beginning

Inventory

+ Production - Sales =

Ending

Inventory

January 700 + 600 - 500 = 800
February 800 + 600 - 250 = 1,150
March 1,150 + 600 - 1,000 = 650